# ToxicTags: Decoding Toxic Memes with Rich Tag Annotations

### v1

@@ Abstract
The 2025 Global Risks Report [^1] https://reports.weforum.org/docs/WEF˙Global˙Risks˙Report˙2025.pdf identifies state-based armed conflict and societal polarisation among the most pressing global threats, with social media playing a central role in amplifying toxic discourse. Memes, as a widely used mode of online communication, often serve as vehicles for spreading harmful content. However, limitations in data accessibility and the high cost of dataset curation hinder the development of robust meme moderation systems. To address this challenge, in this work, we introduce a first-of-its-kind dataset of 6,300 real-world meme-based posts annotated in two stages: (i) binary classification into toxic and normal, and (ii) fine-grained labelling of toxic memes as hateful, dangerous, or offensive. A key feature of this dataset is that it is enriched with auxiliary metadata of socially relevant tags, enhancing the context of each meme. In addition, we propose a tag generation module that produces socially grounded tags, because most in-the-wild memes often do not come with tags. Experimental results show that incorporating these tags substantially enhances the performance of state-of-the-art VLMs in toxicity detection tasks. Our contributions offer a novel and scalable foundation for improved content moderation in multimodal online environments. Warning: Contains potentially toxic contents.
@@ 1 Introduction
Identifying such harmful content is particularly challenging due to the subtle and context-dependent nature of memes. Their meaning is often embedded in cultural references, online trends, sarcasm, or coded language, making them difficult to interpret not only for automated systems but even for human moderators. In response to these challenges, vision language models (VLMs) have recently gained traction as powerful tools for content moderation. These models are capable of jointly analyzing visual and textual elements to grasp the nuanced context of memes and can provide detailed justifications for their classifications (Qu, Backes, and Zhang 2025). Despite these advances, recent research highlights the limitations of VLMs in accurately detecting hateful memes (Rizwan et al. 2025), underscoring the urgent need to bridge these performance gaps. Similarly, a recent blog post [^3] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes/ by Meta highlighted the challenges and complexities inherent in their current content moderation frameworks.
@@
Dataset scarcity: Datasets serve as the foundational fuel for generative AI models. However, researchers increasingly face significant obstacles in curating such datasets, primarily due to the high costs of manual annotation and the limitations imposed on large-scale crawling of social media platforms. These challenges have resulted in datasets that are either manually constructed (Kiela et al. 2020; Lin et al. 2024) – often failing to fully capture the richness of human creativity – or narrowly scoped to specific targets or events (Pramanick et al. 2021a, b; Fersini et al. 2022). Moreover, this scarcity of comprehensive datasets has hindered efforts to simulate the complexity of social media ecosystems across a broad taxonomy of content labels. This is in contrast to textual hate speech research, where more extensive taxonomical explorations exist (Sachdeva et al. 2022).
Our contributions: To address the aforementioned limitations, we make the following contributions.
@@
We introduce a novel, highly diverse, real-world meme dataset enriched with human-annotated contextual tags – a critical yet often overlooked feature in social media content. Unlike prior datasets, it consists of only real-world memes with no restrictions based on specific targets or events (see Section 3). The final dataset comprises 6,300 annotated memes, capturing a wide spectrum of online discourse.
@@
We design a rigorous two-stage human annotation pipeline. In the first stage, memes are classified as either toxic or normal. In the second stage, toxic memes are further categorized into one of three fine-grained classes: hateful, dangerous, or offensive. These categories have been distilled through an iterative annotation process, revealing that a four-class taxonomy is appropriately expressive for moderating a wide range of social media content. This taxonomy can help reduce misclassifications in existing moderation pipelines.
We propose a novel tag generation framework (see Section 6) significantly departing from existing approaches. Leveraging real-world context, our system integrates Google Search API and Google Lens to produce high-quality, socially grounded tags for memes.
We perform few-shot prompting of VLMs using various schemes to select the samples. Our experiments demonstrate that the inclusion of contextual tags significantly enhances model performance in meme classification tasks.
Finally, we report the performance of our method on a standard hate meme detection dataset. Note that this dataset does not have the tag information in the first place; therefore, we use our tag generation module to predict relevant tags for a query image. Inclusion of these tags in the scheme for sampling the few-shot examples results in a better performance compared to the case where the tags are absent.
@@ 2 Related works
Hate meme detection: Rapid surge of hate around the globe (Arcila Calderón et al. 2024) in the recent past years, with memes acting as a major source of fuel, has led to the curation of multiple hateful memes dataset (Kiela et al. 2020; Lin et al. 2025). Similarly, there has been extensive research in the field to build robust content moderation frameworks (Rizwan et al. 2025; Das and Mukherjee 2023b; Prakash et al. 2023; Cao, Lee, and Jiang 2024; Cao et al. 2023), with some works on low-resource languages (Das and Mukherjee 2023a; Kumari et al. 2024). Despite such rapid developments, current datasets are either limited to manually curated memes or focus only on a subset of events (Pramanick et al. 2021a, b; Chen et al. 2023).
dataset
event/target
independent?
real-world
memes?
post’s contextual
information?
two stage
annotation?
size labels
FHM (Kiela et al. 2020) $\sim$10,000 hateful, not-hateful
MAMI (Fersini et al. 2022) $\sim$10,000 misogynistic, not-misogynistic
HARM-C (Pramanick et al. 2021a) $\sim$3,544 very harmful, partially harmful, harmless
HARM-P (Pramanick et al. 2021b) $\sim$3,552 very harmful, partially harmful, harmless
UA-RU Conflict  (Thapa et al. 2022) $\sim$5,680 hateful, not-hateful
CrisisHateMM (Bhandari et al. 2023) $\sim$4,723 hateful, not-hateful
RUHate-MM (Thapa et al. 2024) $\sim$20,675 hateful, not-hateful
ToxicTags (ours) $\sim$6,300
I– toxic, normal
II– hateful, dangerous, offensive, normal
Table 1: Comparison of our dataset with existing datasets on toxic memes. Here post’s contextual information refers to the title or tags associated with the meme (as in ToxicTags).
Tags: Prior research has extensively explored automated tag generation for images (Huang et al. 2024; Zhang et al. 2024; Dai et al. 2023), leading to the development of several dedicated models and datasets. However, to the best of our knowledge, no existing work addresses the specific challenge of tag generation for memes – visual artefacts that often combine text and imagery in culturally nuanced, context-dependent ways. To fill this gap, we curate a dedicated meme-tagging dataset and introduce a novel framework built around KeyLLM [^4] https://maartengr.github.io/KeyBERT/guides/keyllm.html to generate socially relevant and context-aware tags from memes (see Section 6).
This work: We present, for the first time, a diverse real-world dataset of memes labelled across four nuanced categories – toxic, hateful, dangerous, offensive – alongside normal, moving beyond the traditional binary classification used in prior studies (refer to Table 1). In addition, we introduce a novel, socially aware tag generation module, opening new avenues in open-set image tagging. Finally, we employ these tags to improve the performance of toxic meme detection.
3 Dataset
@@
For our dataset, we select three high-volume and thematically rich streams [^6] https://imgflip.com/m/Dark˙humour, https://imgflip.com/m/MEMES˙OVERLOAD, https://imgflip.com/m/politicsDark_Humour, Memes_Overload, and Politics based on their popularity and relevance to diverse meme content. This selection has been made through manual inspection of stream descriptions and content samples (see Appendix C for details). Further, to ensure that the memes included are socially engaging and contextually rich, we manually review a subset of posts and observe that memes with fewer than two comments are often unengaging or irrelevant to a broader audience. Hence, we retain only those memes that received at least two comments. In total, we manually curate 37,072 memes across the selected streams over a period of approximately one month. The memes were downloaded using a browser-based image downloader extension [^7] https://chromewebstore.google.com/detail/download-all-images/ifipmflagepipjokmbdecpmjbibjnakm?hl=en. As a result, our dataset offers a large and diverse collection of real-world memes, free from artificial filtering or event-driven bias.
Metadata: For each post, along with the meme, we collect its (i) title, (ii) number of comments and (iii) the tags list. For the collected memes, we extract the embedded text using GoogleOCR [^8] https://cloud.google.com/vision/docs/ocr. To ensure robustness of OCR extracted text, we perform manual verification of 100 randomly chosen samples and find nearly 98 to be correctly identified; thus depicting the robustness of the tool. Further, we also store the bounding box of OCR and use it later for generating tags (refer to Section 6).
@@
Agreement: We strictly adhere to the privacy and copyright regulations of the platform [^10] https://imgflip.com/terms and therefore collected our data only from the publicly available posts. To maintain the privacy of users, we did not store any information that could potentially compromise their anonymity.
@@ 4 Annotation
Two staged annotation: We adopt a two stage annotation process – (i) classification of each meme into toxic or normal, and (ii) further bifurcation of toxic memes into offensive, dangerous or hateful. The two-stage process is specifically adopted to mitigate annotation errors as suggested in (Rizwan et al. 2025). Further, we also conduct a pilot study at each stage before concluding the final annotations. For both stages, we perform three annotations per sample by three different annotators to minimise annotation bias, given that this is a highly subjective task (Mathew et al. 2022). We use the standard definition of labels that are effectively used for practical applications. In the upcoming subsections, we discuss each stage separately and further present detailed annotation instructions with corresponding definitions of the labels in Appendix D (also see Figure 5 in the Appendix).
Annotator details: We initially selected 25 annotators through an open call in a university setting. The minimum criteria that we set are (i) the enrollment of the candidate in the second year of their bachelor’s program, and (ii) familiarity of the candidate with diverse social media content. Among these, five are experienced researchers having exposure to this field, and the rest 20 candidates are engineering students in their sophomore years. All the annotators are within the age range of 21-27 years, and the whole annotation process has been conducted under the close supervision of two senior researchers. To help the annotators with their regular queries and to make sure their mental health is not affected, we conducted daily scrum and further made a Slack workspace for easier collaboration. In particular, we asked all the participating annotators to sign an agreement before starting the annotation, wherein we specifically focus on their mental well-being, given the nature of the work. We detail out the safety guidelines in Appendix D; further, as an extra safety measure, we rolled out a maximum of only 50 samples per day to each annotator.
Annotation tool: Separate PDFs for memes with all required metadata (as shown in Figure 5 in the Appendix) and a separate Google sheet have been provided to each annotator to reduce bias and ensure fair annotation. Further in our scrum with the annotators, we assisted them by discussing their queries to provide them further insights so as to maintain a high quality of annotation.
@@
Annotation codebook: We also provide detailed annotation instructions, accompanied by 25 samples that have been manually selected and annotated by the two supervising researchers. As outlined in Appendix D, our annotation codebook has been constructed based on standard guidelines from which we have derived our definitions. For the toxic label, we strictly follow the definition used by the Perspective API, given its widespread adoption in production settings. For the hateful label, we adhere to the Facebook hate meme definition (Kiela et al. 2020), which has become the de facto standard. For the dangerous label, we follow the annotation guidelines available at the following source [^11] https://www.dangerousspeech.org/dangerous-speech.
Stage I: Binary labelling of the dataset
@@
Annotation: The full set of 6,300 samples has been evenly distributed among the 12 selected annotators (out of the 25 recruited). Each post has been independently annotated by three annotators. We achieve a final inter-annotator agreement score of 0.8176, as measured by Fleiss’ $\kappa$. Notably, for a subjective task (Mathew et al. 2022) such as ours, this level of agreement is considered relatively high, thereby validating the quality of our annotations. Each annotator has been paid USD 40 for this task which is much above the minimum wage in the annotators’ country.
@@
Label assignment: The final label for each sample is determined based on the majority vote of the three annotations, i.e., the label with at least two agreements is considered to be the final label. In this stage, we finally obtained 1,392 normal and 4908 toxic samples (refer to Table 2 for complete statistics).
Stage II: Fine-grained labelling
@@
Annotation: As a pilot step, we randomly sampled 250 toxic memes from the previous stage and had the expert annotators categorise them into three distinct labels. These labelled samples were then evenly distributed among the 12 annotators, each of whom annotated all the assigned samples. Upon verification, we found that the annotators were performing satisfactorily, and thus proceeded with the final annotation with this group. Subsequently, the set of 4,908 toxic samples was categorised into three classes: hateful, dangerous, and offensive. As before, each sample received three independent annotations. The annotators achieved a Fleiss’ $\kappa$ agreement score of 0.8197 – a slight improvement over Stage I – highlighting the value of multi-stage and multi-pilot studies in enhancing annotation quality.
@@
Label assignment: The final label is determined based on the majority vote of the annotators for each sample. When no label receives more than one vote (which means equal vote to hateful, dangerous and offensive for that particular sample), we exclude that sample from the dataset and put it in undecided category. The overall dataset statistics are noted in Table 2. To maintain the diversity of tags in both the splits, we ensure that each tag appears at least 15% of the times in the test split, resulting in a total of 1000 test samples and 5,300 training samples. Word clouds for label wise tags are presented in Table 9 of Appendix E.
@@
I & II normal 1243 149 1392
I toxic 4057 851 4908
hateful 1475 343 1818
dangerous 1788 375 2163
offensive 653 122 775
II undecided 141 11 152
@@
Table 2: Data statistics showing the distribution of the data points across the classes.

### v2

# StemTox: From Social Tags to Fine-Grained Toxic Meme Detection via Entropy-Guided Multi-Task Learning

@@ Abstract
Memes, as a widely used mode of online communication, often serve as vehicles for spreading harmful content. However, limitations in data accessibility and the high costs of dataset curation hinder the development of robust meme moderation systems. To address this challenge, in this work, we introduce a first-of-its-kind dataset – ToxicTags consisting of 6,300 real-world meme-based posts annotated in two stages: (i) binary classification into toxic and normal, and (ii) fine-grained labelling of toxic memes as hateful, dangerous, or offensive. A key feature of this dataset is that it is enriched with auxiliary metadata of socially relevant tags, enhancing the context of each meme. In addition, we propose a novel entropy guided multi-tasking framework – StemTox – that integrates the generation of socially grounded tags with a robust classification framework. Experimental results show that incorporating these tags substantially enhances the performance of state-of-the-art VLMs in toxicity detection tasks. Our contributions offer a novel and scalable foundation for improved content moderation in multimodal online environments. Warning: Contains potentially toxic contents.
@@ 1 Introduction
Identifying such harmful content is particularly challenging due to the subtle and context-dependent nature of memes. Their meaning is often embedded in cultural references, online trends, sarcasm, or coded language, making them difficult to interpret not only for automated systems but even for human moderators. In response to these challenges, vision language models (VLMs) have recently gained traction as powerful tools for content moderation. These models are capable of jointly analyzing visual and textual elements to grasp the nuanced context of memes and can provide detailed justifications for their classifications Qu et al. (2025). Despite these advances, recent studies have highlighted the limitations of VLMs in accurately detecting hateful memes Rizwan et al. (2025), underscoring the urgent need to bridge these performance gaps. Similarly, a recent blog post [^2] https://about.fb.com/news/2025/01/meta-more-speech-fewer-mistakes/ by Meta highlighted the challenges and complexities inherent in their current content moderation frameworks.
@@
Datasets serve as the foundational fuel for generative AI models. However, researchers increasingly face significant obstacles in curating such datasets, primarily due to the high costs of manual annotation and the limitations imposed on large-scale crawling of social media platforms. These challenges have resulted in datasets that are either manually constructed Kiela et al. (2020); Lin et al. (2024) – often failing to fully capture the richness of human creativity – or narrowly scoped to specific targets or events Pramanick et al. (2021a, b); Fersini et al. (2022). Moreover, this scarcity of comprehensive datasets has hindered efforts to build robust toxicity detection systems for multimodal social media content. This is in contrast to textual hate speech research, where more extensive taxonomical explorations exist Sachdeva et al. (2022).
In this work, we present, for the first time, a diverse real-world dataset of memes labelled across four nuanced categories – toxic, hateful, dangerous, offensive – alongside normal, moving beyond the traditional binary classification paradigm predominantly used in prior studies (refer to Table 1). In addition, each meme in the dataset is also labelled with a set of social tags that contextually ground the meme. Finally, we introduce StemTox, a novel, socially aware entropy-guided multi-tasking framework, opening new avenues for open-set image tagging and robust toxicity classification. Our contributions are noted below.
@@
The key contributions of our work are as follows: ☞ We introduce a novel, highly diverse, real-world meme dataset ToxicTags enriched with human-annotated contextual tags – a critical yet often overlooked feature in social media content. Unlike prior datasets, it consists of only real-world memes with no restrictions based on specific targets or events (see Section 3). The final dataset comprises 6,300 annotated memes, capturing a wide spectrum of online discourse.
@@
☞ We design a rigorous two-stage human annotation pipeline. In the first stage, memes are annotated as either toxic or normal. In the second stage, toxic memes are further categorized into one of three fine-grained classes: hateful, dangerous, or offensive. These categories have been distilled through an iterative annotation process, revealing that a four-class taxonomy is appropriately expressive for moderating a wide range of social media content. This taxonomy can help reduce misclassifications in existing toxicity detection pipelines.
☞ We introduce StemTox, a novel framework that significantly advances the state-of-the-art by generating socially grounded, high-quality tags and improving meme classification performance. Our experiments demonstrate that incorporating contextual tags into the multitasking system leads to significant performance gains in meme classification tasks.
☞ Finally, we report the performance of our method on two standard benchmark hate meme detection datasets and compare it against five baseline approaches.
@@ 2 Related Works
Hate meme detection: Rapid surge of hate around the globe Arcila Calderón et al. (2024) in the recent past years, with memes acting as a major source of fuel, has led to the curation of multiple hateful memes dataset Kiela et al. (2020); Lin et al. (2025). Similarly, there has been extensive research in the field to build robust content moderation frameworks Rizwan et al. (2025); Das and Mukherjee (2023b); Prakash et al. (2023); Cao et al. (2024a, 2023b), with some works on low-resource languages Das and Mukherjee (2023a); Kumari et al. (2024). Despite such rapid developments, current datasets are either limited to manually curated memes or focus only on a subset of events Pramanick et al. (2021a, b); Chen et al. (2023).
dataset event / target independent? real world memes? post’s contextual information? two stage annotation? size labels
FHM Kiela et al. (2020) $\sim$10k hateful, not-hateful
MAMI Fersini et al. (2022) $\sim$10k misogynistic, not-misogynistic
HARM-C Pramanick et al. (2021a) $\sim$3,544 very/partially harmful, harmless
HARM-P Pramanick et al. (2021b) $\sim$3,552 very/partially harmful, harmless
UA–RU Conflict Thapa et al. (2022) $\sim$5,680 hateful, not-hateful
CrisisHateMM Bhandari et al. (2023) $\sim$4,723 hateful, not-hateful
RUHate-MM Thapa et al. (2024) $\sim$20,675 hateful, not-hateful
ToxicTags $\sim$6,300 stage I: toxic, normal stage II: hateful, dangerous, offensive, normal
Table 1: Comparison of ToxicTags with existing datasets for toxic meme detection. Post-contextual information includes metadata such as titles or tags associated with the meme.
Tags: Prior research has extensively explored automated tag generation for images Huang et al. (2024); Zhang et al. (2024); Dai et al. (2023), leading to the development of several dedicated models and datasets. However, to the best of our knowledge, no existing work addresses the specific challenge of tag generation for memes – visual artefacts that often combine text and imagery in culturally nuanced, context-dependent ways.
Present work: We present a real-world meme dataset with fine-grained toxicity labels beyond binary classification (Table 1). In addition, each meme is annotated with a set of socially aware tags. We also introduce a StemTox for open-set image tagging and robust classification.
3 The ToxicTags dataset
@@
For our dataset, we select three high-volume and thematically rich streams—Dark_Humour [^4] https://imgflip.com/m/Dark_humour, Memes_Overload [^5] https://imgflip.com/m/MEMES_OVERLOAD, and Politics [^6] https://imgflip.com/m/politics—based on their popularity and relevance to diverse meme content. We selected these streams through manual inspection of their descriptions and content samples(see Appendix C for details). Further, to ensure that the memes included are socially engaging and contextually rich, we manually review a subset of posts and observe that memes with fewer than two comments are often unengaging or irrelevant to a broader audience. Hence, we retain only those memes that received at least two comments. In total, we manually curated 37,072 memes across the selected streams over a period of approximately one month. The memes were downloaded using a browser-based image downloader extension [^7] https://chromewebstore.google.com/detail/download-all-images/ifipmflagepipjokmbdecpmjbibjnakm?hl=en. As a result, our dataset offers a large and diverse collection of real-world memes, free from artificial filtering or event-driven bias.
Metadata: For each post, along with the meme, we collect its (i) title, (ii) number of comments and (iii) the tags list. For the collected memes, we extract the embedded text using GoogleOCR [^8] https://cloud.google.com/vision/docs/ocr. To ensure robustness of OCR extracted text, we perform manual verification of 100 randomly chosen samples and find nearly 98 to be correctly identified; thus depicting the robustness of the tool.
@@
Agreement: We strictly adhered to the privacy and copyright regulations of the platform [^10] https://imgflip.com/terms and therefore collected our data only from the publicly available posts. To maintain the privacy of users, we did not store any information that could potentially compromise their anonymity.
@@ 4 Annotation
Two staged annotation: We adopt a two stage annotation process – (i) classification of each meme into toxic or normal, and (ii) further bifurcation of toxic memes into offensive, dangerous or hateful. The two-stage process is specifically adopted to mitigate annotation errors as suggested in Rizwan et al. (2025).
@@
Figure 1 presents a brief pipeline of the employed annotation process over two stages. The instructions that we designed are given below, and some samples from those provided to annotators are present in Table 10. During the whole annotation process, annotators were asked to strictly adhere to the label definitions (refer Appendix D). Further, we also conducted a pilot study at each stage before progressing to the final annotations. For both stages, we perform three annotations per sample by three different annotators to minimise annotation bias, given that this is a highly subjective task Mathew et al. (2022). We use the standard definition of labels that are effectively used for practical applications. In the upcoming subsections, we discuss each stage separately and further present detailed annotation instructions with corresponding definitions of the labels in Appendix D.
Annotator details: To ensure a high-quality annotation process, we shortlisted 25 annotators based on the following minimum criteria: (i) candidates having professional experience and prior work aligned with our research, and (ii) familiarity with diverse social media content. We additionally required annotators to pass a qualification task curated by the supervising researchers to assess label comprehension and consistency. The entire annotation process was conducted under the close supervision of two PhD researchers. We detail the safety guidelines in Appendix D; further, as an extra safety measure, we rolled out a maximum of only 50 samples per day to each annotator.
@@
Annotation codebook: We also provide detailed annotation instructions, accompanied by 25 samples that have been manually selected and annotated by the two supervising researchers. As outlined in Appendix D, our annotation codebook has been constructed based on standard guidelines from which we have derived our definitions. For the toxic label, we strictly follow the definition used by the Perspective API, given its widespread adoption in production settings. For the hateful label, we adhere to the Facebook hate meme definition Kiela et al. (2020), which has become the de facto standard. For the dangerous label, we follow the annotation guidelines available at the following source  [^11] https://www.dangerousspeech.org/dangerous-speech.
4.1 Stage I: Binary labelling of the dataset
@@
Annotation: The full set of 6,300 samples has been evenly distributed among the 12 selected annotators (out of the 25 recruited). As stated earlier, we employ three annotators per sample which enhances annotation diversity. We achieve a final inter-annotator agreement score of 0.753, as measured by Fleiss’ $\kappa$. Notably, for a subjective task Mathew et al. (2022) such as ours, this level of agreement is considered relatively high, thereby validating the quality of our annotations. Each annotator has been paid USD 40 for this task which is much above the minimum wage in the annotators’ country.
@@
Label assignment: The final label for each sample is determined based on the majority vote of the three annotations, i.e., the label with at least two agreements is considered to be the final label. In this stage, we finally obtained 1,446 normal and 4,854 toxic samples (refer to Table 2 for complete statistics).
4.2 Stage II: Fine-grained labelling
@@
Annotation: As a pilot step, we randomly sampled 250 toxic memes from the previous stage and had the expert annotators categorise them into three distinct labels. These labelled samples were then evenly distributed among the 12 annotators, each of whom annotated all the assigned samples. Upon verification, we found that the annotators were performing satisfactorily, and thus proceeded with the final annotation with this group. Subsequently, the set of 4854 toxic samples was categorised into three classes: hateful, dangerous, and offensive. As before, each sample received three independent annotations. The annotators achieved a Fleiss’ $\kappa$ agreement score of 0.793 – a slight improvement over Stage I – highlighting the value of multi-stage and multi-pilot studies in enhancing annotation quality.
@@
Label assignment: The final label is determined based on the majority vote of the annotators for each sample. When no label receives more than one vote (which means equal vote to hateful, dangerous and offensive for that particular sample). For undecidable cases, final annotation was carried out by our expert researchers. The overall dataset statistics are noted in Table 2. To maintain the diversity of tags in both the splits, we ensure that each tag appears at least 15% of the times in the test split, resulting in a total of 1000 test samples and 5,300 training samples. Word clouds for label wise tags are presented in Table 11 of Appendix E.
@@
I & II normal 1297 149 1446
I toxic 4003 851 4854
II hateful 1470 282 1752
dangerous 1847 472 2319
offensive 686 97 783
@@
Table 2: Dataset statistics showing the distribution of samples across classes.
tag pairs theme
Table 3: Top co-occurring tag pairs and associated semantic themes.
Figure 3: Proposed StemTox framework.